OpenStudy (anonymous):

4t+7-t=19 HELP MEEEE

OpenStudy (anonymous):

4t-t=19-7 so you bring the 7 over and make it a negative and because they are like terms! so it would equal 3t=12 and divide 3 on both sides so 3t/3=12/3, t=4 ! hope this helped :)